Nonconstant Growth. Better Mousetraps has come out with an improved product, and the world is beating a path to its door. As a result, the firm projects growth of 20% per year for 4 years. By then, other firms will have copycat technology, competition will drive down profit margins, and the sustainable growth rate will fall to 5%. The most recent annual dividend was DIV0 = $1 per share. (LO7-2)

a. What are the expected values of: (i) DIV1, (ii) DIV2, (iii) DIV3, and (iv) DIV4?

b. What is the expected stock price 4 years from now? The discount rate is 10%.

c. What is the stock price today?
Answer #1

a). D1 = D0 * (1 + g) = $1 * (1 + 0.20) = $1.20

D2 = D1 * (1 + g) = $1.20 * (1 + 0.20) = $1.44

D3 = D2 * (1 + g) = $1.44 * (1 + 0.20) = $1.728

D4 = D3 * (1 + g) = $1.728 * (1 + 0.20) = $2.0736

b). P4 = [D4 * (1 + gC)] / [r - gC]

= [$2.0736 * (1 + 0.05)] / [0.10 - 0.05] = $2.17728 / 0.05 = $43.5456, or $43.55

c). P0 = [D1 / (1 + r)] + [D2 / (1 + r)2] + [D3 / (1 + r)3] + [(D4 + P4) / (1 + r)4]

= [$1.20 / (1 + 0.10)] + [$1.44 / (1 + 0.10)2] + [$1.728 / (1 + 0.10)3] + [($2.0736 + $43.55) / (1 + 0.10)4]

= $1.09 + $1.19 + $1.30 + $31.16 = $34.74